40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 25 17:42:11 2021, Anchal, Paco, Summary, ALS, ALS Single Arm Noise Budget ALS_IR_b.svgALS_Single_Arm_IR.pdf
    Reply  Thu May 27 11:03:15 2021, Anchal, Paco, Summary, ALS, ALS Single Arm Noise Budget ALS_Single_X_Arm_IR.pdfALS_OOL_with_Ref.pdf
       Reply  Fri May 28 17:32:48 2021, Anchal, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat SingleArmActCalwithIRALSBeat.pdfstateSpaceModel.zip
          Reply  Tue Jun 1 16:55:32 2021, Anchal, Paco, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat SingleArmActCalwithIRALSBeat-1306624785.pdf
             Reply  Tue Jun 8 11:40:53 2021, Anchal, Paco, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat ITMX_Cal_Noise_Spectrum_1307143423.pdf
                Reply  Wed Jun 9 18:29:13 2021, Anchal, Paco, Summary, ALS, Check for saturation in ITMX SOS Driver 
                Reply  Fri Jul 9 15:39:08 2021, Anchal, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat [Correction] ITMX_calibration_With_ALS_Beat.pdf
Message ID: 16171     Entry time: Tue Jun 1 16:55:32 2021     In reply to: 16168     Reply to this: 16192
Author: Anchal, Paco 
Type: Summary 
Category: ALS 
Subject: Single Arm Actuation Calibration with IR ALS Beat 

Rana suggested in today's meeting to put in a notch filter in the XARM IR PDH loop to avoid suppressing the excitation line. We tried this today first with just one notch at 1069 Hz and then with an additional notch at 619 Hz and sent two simultaneous excitations.


Measurement and Analysis:

  • We added notch filters with Q=10, depth=50dB, freq=619 Hz and 1069 Hz using foton in SUS-ETMX_LSC filter bank at FM10.
  • We sent excitation signals with amplitudes 600cts and 1000 cts for 619 Hz and 1069 Hz signals respectively.
  • We measured time series data of C1:SUS-ITMX_LSC_OUT_DQ and C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ for 60s.
  • Then, spectrum of both signals is measured with Hanning window using scipy.welch function with scaling set to  'spectrum', binwidth=1Hz.
  • The beatnote signal was converted into length units by multiplying it by 1064nm * 37.79m / c.
  • The ratio of the two spectrums at teh excitation frequency multiplies by excitation frequency squared gives us teh calibration constant in units of nm Hz^2/cts.
  • At 619 Hz, we got \frac{5.01}{f^2}nm/cts
  • At 1069 Hz, we got \frac{5.64}{f^2}nm/cts.
  • The calibration factor in use is from \frac{7.32}{f^2} nm/cts from 13984.
  • So, the calibration factor from this methos is about 23% smaller than measured using freeswinging MICH in 13984.
  • One possiblity is that our notch filter is not as effective in avoiding suppresion of excitation.
    • We tried increasing the notch filter depths to 100 dB but got the same result within 2%.
    • We tried changing the position of notch filters. We put them in POX filter banks. Again the result did not change more than 2%.
  • The open loop gain of green PDH at 619 Hz and 1069 Hz must be large enough for our assumption of green laser perfectly following length motion to be true. The UGF of green laser is near 11 kHz.
  • The discrepancy could be due to outdated freeswinging MICH measurement that was done 3 years ago. Maybe we should learn how to do the ITMX calibration using this method and compare our own two measurements.
Attachment 1: SingleArmActCalwithIRALSBeat-1306624785.pdf  26 kB  Uploaded Tue Jun 1 17:57:07 2021  | Show | Show all
ELOG V3.1.3-